Most of the wind associated with a thunderstorm is caused by parcels of air, called air masses, that fall down due to gravity (which is called a down draft). a) Assuming that the air mass falls down from the top of a thunderstorm which is 12 km above the ground, how fast is the air mass traveling when it reaches the ground? b) Is this speed reasonable
1
Expert's answer
2019-05-21T08:50:42-0400

a) We have:


"\rho gh=0.5\rho v^2"

"v=\sqrt{2gh}=\sqrt{2(9.8)(12000)}=485\frac{m}{s}"

b) No.